10 out of 10 points Question 1 A population has a mean of 60 and a standard deviation of 30. Samples of size 16 are randomly Please help with 5,6,7, please show your work as well!
Answer #1

3 и – чоооо 6 - 8 боо n = 400 к (ху чо5оо ) - e(Х-и s > 405oo чоооо 800 - (> 1.25) 6. JOS6 y: ts – goo, Б- So ост до (844 2 8n=64 X=so, s=5, tool ş =t63, o os 2.00 2 me= tn-l m - 2 x vra 1-25 lower limit = ame =50 -1.25 = 48.75 Upper limit=7+me -30 1